# spacing=nonfrench

### 04 Mar 2026 v1

@@ 2.2. Fontaine’s period rings
Fontaine achieved a classification of the $p$-adic representations of the absolute Galois group $G_{E}$ of a finite extension $E/\mathbb{Q}_{p}$. His classifier relies on rings such as $B_{\mathrm{dR}}$ and $B_{\mathrm{cris}}$. In this subsection, we outline the construction of these two rings and review the formal group setting. The original constructions can be found in [Fon94] and in [Col02] for the tensorial generalization. To coincide with our situation, we work directly in the particular case of $E=L$.
@@ 2.2.1. The integers of the tilt of $\mathbb{C}_{p}$
Let $\mathcal{O}_{\mathbb{C}_{p}}\!$ be the closed unit disk of $\mathbb{C}_{p}$. Throughout the remainder of this article, we fix $\omega_{L}$ a uniformizer of $L$ and let $q$ be the cardinality of the residue field $\mathcal{O}_{L}/\mathfrak{m}_{L}=\mathbb{F}_{q}$. The definition of the aforementioned period rings involves several mechanisms, whose starting point is the perfect ring
@@
of characteristic $p$. It is endowed with a valuation $\upsilon_{\mathcal{R}}$ for which it is complete, defined by $\upsilon_{\mathcal{R}}(x)=\lim_{n\to+\infty}q^{n}\upsilon_{p}(\widehat{x}_{n})$ for all $x=(x_{n})_{n\geqslant 0}\in\mathcal{R}$, where $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}}\!$ is a lift of $x_{n}$ modulo $(\omega_{L})$ for all integers $n\geqslant 0$ and $\upsilon_{p}$ is the $p$-adic valuation normalized on $\mathbb{Q}_{p}$. This ring is also equipped with a compatible action of $G_{L}$ inherited from its action on $\mathcal{O}_{\mathbb{C}_{p}\!}$.
@@ 2.2.2. Blank period rings
We now pass to characteristic zero. Let $\operatorname{W}_{\!\omega_{L}\!}$ be the generalized $\omega_{L}$-adic Witt vectors endofunctor of $\mathcal{O}_{L}$-algebras; see [CD15, §2] for a presentation. The ring $\mathcal{R}$ is naturally an $\mathcal{O}_{L}$-algebra and we let
$A_{\mathrm{inf},L}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{W}_{\!\omega_{L}\!}(\mathcal{R})$
be the unique strict $\omega_{L}$-ring of residue ring $\mathcal{R}$. Alternatively, one can see this blank period ring as $\mathcal{O}_{L}\otimes_{\operatorname{W}(\mathbb{F}_{q})}\operatorname{W}(\mathcal{R})$, where $\operatorname{W}$ is the usual $p$-adic Witt vectors functor. Elements of this ring can be uniquely written in the form $\sum_{n\geqslant 0}[x_{n}]\omega_{L}^{n}$ where $x_{n}\in\mathcal{R}$ and [][-] is the Teichmüller lift. The action of $G_{L}$ on $\mathcal{R}$ lifts by functoriality to an action on $A_{\mathrm{inf},L}$. An important map attached to it is the $G_{L}$-equivariant epimorphism $\theta\colon A_{\mathrm{inf},L}\twoheadrightarrow\mathcal{O}_{\mathbb{C}_{p}}\!$ of $\mathcal{O}_{L}$-algebras uniquely determined by mapping $[x]$ to $\lim_{n\to+\infty}\widehat{x}_{n}^{q^{n}}\!$, where $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}}\!$ is a lift of the $n$-th component of $x\in\mathcal{R}$. This map extends in a unique way by $L$-linearity to a $G_{L}$-equivariant epimorphism $\theta$ from $B_{\mathrm{inf},L}^{+}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=A_{\mathrm{inf},L}[1/\omega_{L}]$ to $\mathbb{C}_{p}$, whose kernel is principal.
@@ 2.2.3. The de Rham period field
is the $\ker(\theta)$-adic completion of $B_{\mathrm{inf},L}^{+}$. It inherits the topology generated by the intersection of the $\ker(\theta)$-adic and $\omega_{L}$-adic topologies from its dense subring of definition, along with a compatible action of $G_{L}$. Once again, $\theta$ extends to a $G_{L}$-equivariant epimorphism $\theta\colon B_{\mathrm{dR}}^{+}\twoheadrightarrow\mathbb{C}_{p}$ and its kernel is still principal. The $\mathbb{Z}$-powers of $\ker(\theta)$ induce a decreasing, exhaustive and separated filtration $(\operatorname{Fil}^{n}B_{\mathrm{dR}})_{n\in\mathbb{Z}}$ on the fraction field $B_{\mathrm{dR}}$ of $B_{\mathrm{dR}}^{+}$. This field is our first ring of interest.
The field $B_{\mathrm{dR}}$ is sufficient to capture $p$-adic representations arising from formal groups, but too coarse to be used to classify them. Denote by $\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}(G_{L})$ the category of $p$-adic representations of $G_{L}$, and by $\operatorname{\mathbf{Fil}}_{L}$ the one of finite-dimensional $L$-vector spaces equipped with a decreasing, exhaustive and separated $\mathbb{Z}$-filtration by subspaces. Fontaine defined a contravariant functor
@@
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{dR}})$, equipped with the induced filtration $\operatorname{Fil}^{n}\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,\operatorname{Fil}^{n}B_{\mathrm{dR}})$ for all $n\in\mathbb{Z}$. We say that a $p$-adic representation $V$ of $G_{L}$ is de Rham if the $L$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. In this case, the jumps of the filtration are called the Hodge-Tate weights of $V$. Restricting the preceding functor to de Rham representations yields a faithful functor. When $(V,\rho)$ is one-dimensional, being de Rham is equivalent to the existence of a nonzero element $t\in B_{\mathrm{dR}}$ such that $g.t=\rho(g)t$ for all $g\in G_{L}$. Then, the Hodge-Tate weight is the largest $w\in\mathbb{Z}$ such that $t\in\operatorname{Fil}^{w}B_{\mathrm{dR}}$.
@@ 2.2.4. The crystalline period ring
To refine the theory, Fontaine defined a subring $B_{\mathrm{cris}}$ of $B_{\mathrm{dR}}$ as follows: let $A_{\mathrm{cris}}^{\circ}$ be the PD-envelope of $\operatorname{W}(\mathcal{R})$ with respect to $\ker(\theta)$. Set $A_{\mathrm{cris}}$ to be its $p$-adic completion. One can show that the latter is still a subring of $B_{\mathrm{dR}}$. Define $B_{\mathrm{cris}}^{+}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=A_{\mathrm{cris}}[1/p]$. The $G_{L}$-action on $\operatorname{W}(\mathcal{R})$ extends at each previous level and the latter contains a unique $G_{L}$-stable $\mathbb{Q}_{p}$-line on which $G_{L}$ acts by the cyclotomic character. The crystalline period ring
$B_{\mathrm{cris}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=B_{\mathrm{cris}}^{+}[1/t]$
is obtained from $B_{\mathrm{cris}}^{+}$ by inverting any nonzero $t$ in this line. Unlike $B_{\mathrm{dR}}$, the Frobenius $\varphi$ on $\operatorname{W}(\mathcal{R})$ extends uniquely to $B_{\mathrm{cris}}$ by setting $\varphi(t)=pt$, thereby yielding an injective $G_{L}$-equivariant endomorphism $\varphi\colon B_{\mathrm{cris}}\hookrightarrow B_{\mathrm{cris}}$ which is semilinear with respect to the Frobenius $\sigma$ of $L_{0}$, the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$. The canonical map $L\otimes_{L_{0}}B_{\mathrm{cris}}\rightarrow B_{\mathrm{dR}}$ is injective.
The ring $B_{\mathrm{cris}}$ will be of primary interest to us, as it is rich enough not only to capture but also to classify representations arising from formal groups. Let $\operatorname{\mathbf{MF}}_{L}(\varphi)$ denote the category of finite-dimensional $L_{0}$-vector spaces equipped with a bijective $\sigma$-semilinear Frobenius endomorphism $\varphi$, whose scalar extension to $L$ is an object of $\operatorname{\mathbf{Fil}}_{L}$. Morphisms in this category are $L_{0}$-linear maps commuting with $\varphi$ and preserving the filtration after scalar extension. Fontaine defined a contravariant functor
@@
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{cris}})$, equipped with the Frobenius induced by that on $B_{\mathrm{cris}}$. Moreover, one has $L\otimes_{L_{0}}\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)\simeq\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$. We say that a $p$-adic representation $V$ of $G_{L}$ is crystalline if the $L_{0}$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. In this case, its Hodge-Tate weights are its weights as a de Rham representation. In this setting, the previous functor yields an antiequivalence of categories
@@ 2.2.5. Application to formal groups
Given a formal group $H$ of height $h\geqslant 1$ over $\mathcal{O}_{L}$, the $\mathbb{Q}_{p}$-scalar extension of its $p$-adic Tate module
$\operatorname{V}_{\!p}(H)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\operatorname{T}_{p}(H)$
is a de Rham $G_{L}$-representation of dimension $h$ and of Hodge-Tate weight $1$. It is even a crystalline representation of $G_{L}$. In this paper, we are primarily interested in the integral setting and the situation is as follows. Let $\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$ denote the category of formal groups of finite height defined over $\mathcal{O}_{L}$, and $\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$ the category of $G_{L}$-stable $\mathbb{Z}_{p}$-lattices in crystalline representations of $G_{L}$ of Hodge-Tate weights $1$. By [SW13, Corollary 6.2.3] and [Tat67, Proposition 1], the functor
$\operatorname{T}_{p}\colon\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}\longrightarrow\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$
is an equivalence of categories. When $H$ is of height $1$, any endomorphism $s\in\operatorname{End}_{\mathcal{O}_{L}}(H)$ acts on $\operatorname{T}_{p}(H)$ by multiplication by its linear coefficient $s^{\prime}(0)\in\mathbb{Z}_{p}$, and hence extends to an endomorphism of $\operatorname{V}_{\!p}(H)$. In view of the equivalence and the fact that this scalar extension is faithful, the single coefficient $s^{\prime}(0)$ together with the $\mathbb{Z}_{p}$-module structure on $\operatorname{T}_{p}(H)$ or $\operatorname{V}_{\!p}(H)$ uniquely determine the endomorphism $s$. It also acts on $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(\operatorname{V}_{\!p}(H))$ in the following way: for a nonzero $v\in\operatorname{V}_{\!p}(H)$ and a nonzero period $t\in B_{\mathrm{dR}}$ of $\operatorname{V}_{\!p}(H)$, the map induced by $s$ sends $(v\mapsto t)$ to $(v\mapsto s^{\prime}(0)t)$ since $s(v)=s^{\prime}(0)v$. The initial series $s$ is unequivocally determined by this action.

### 22 Apr 2026 v2

@@ 2.2. Fontaine’s period rings
Fontaine has classified the $p$-adic representations of the absolute Galois group $G_{E}$ of a finite extension $E/\mathbb{Q}_{p}$. His classifier relies on period rings such as $B_{\mathrm{dR}}$ and $B_{\mathrm{cris}}$. In this subsection, we outline the construction of these two rings and review the formal group setting. The original constructions can be found in [Fon94], and in [Col02] for the tensorial generalization. To coincide with our situation, we work directly in the particular case of $E=L$.
@@ 2.2.1. The integers of the tilt of $\mathbb{C}_{p}$
Let $\mathcal{O}_{\mathbb{C}_{p}}\!$ be the closed unit disk of $\mathbb{C}_{p}$. Throughout the remainder of this article, we fix $\pi_{L}$ a uniformizer of $L$ and we let $q$ be the cardinality of the residue field $\mathcal{O}_{L}/\mathfrak{m}_{L}=\mathbb{F}_{q}$. The definition of the aforementioned period rings involves several mechanisms, whose starting point is the perfect ring
@@
of characteristic $p$. This ring is endowed with a valuation $\upsilon_{\mathcal{R}}$ for which it is complete, defined by $\upsilon_{\mathcal{R}}(x)=\lim_{n\to+\infty}q^{n}\upsilon_{p}(\widehat{x}_{n})$ for all $x=(x_{n})_{n\geqslant 0}\in\mathcal{R}$, where $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}\!}$ is a lift of $x_{n}$ modulo $(\pi_{L})$ for all integers $n\geqslant 0$ and where $\upsilon_{p}$ is the $p$-adic valuation normalized on $\mathbb{Q}_{p}$. Moreover, this ring is equipped with a compatible action of $G_{L}$ inherited from the action of the latter on $\mathcal{O}_{\mathbb{C}_{p}\!}$.
@@ 2.2.2. Blank period rings
We move to characteristic zero. Let $\operatorname{W}_{\!\pi_{L}\!}$ be the generalized $\pi_{L}$-adic Witt vectors endofunctor of $\mathcal{O}_{L}$-algebras; see for instance [CD15, §2] for a short presentation in our context. The ring $\mathcal{R}$ is naturally an $\mathcal{O}_{L}$-algebra and we let
$A_{\mathrm{inf},L}=\operatorname{W}_{\!\pi_{L}\!}(\mathcal{R})$
be the unique strict $\pi_{L}$-ring of residue ring $\mathcal{R}$. Alternatively, one can see this blank period ring as $\mathcal{O}_{L}\otimes_{\operatorname{W}(\mathbb{F}_{q})}\operatorname{W}(\mathcal{R})$, where $\operatorname{W}$ is the usual $p$-adic Witt vectors functor. Its elements can be uniquely written in the form $\sum_{n\geqslant 0}[x_{n}]\pi_{L}^{n}$, where $x_{n}\in\mathcal{R}$ and [][-] is the Teichmüller lift. The action of $G_{L}$ on $\mathcal{R}$ lifts by functoriality to an action on $A_{\mathrm{inf},L}$. An important map attached to this ring is the $G_{L}$-equivariant epimorphism $\theta\colon A_{\mathrm{inf},L}\twoheadrightarrow\mathcal{O}_{\mathbb{C}_{p}}\!$ of $\mathcal{O}_{L}$-algebras determined by mapping $[x]$ to $\lim_{n\to+\infty}\widehat{x}_{n}^{q^{n}}\!$, where the element $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}}\!$ lifts the $n$-th component of $x\in\mathcal{R}$. This map uniquely extends by $L$-linearity to a $G_{L}$-equivariant epimorphism $\theta$ from $B_{\mathrm{inf},L}^{+}=A_{\mathrm{inf},L}[1/\pi_{L}]$ to $\mathbb{C}_{p}$ and its kernel is principal.
@@ 2.2.3. The de Rham period field
is the $\ker(\theta)$-adic completion of $B_{\mathrm{inf},L}^{+}$. It inherits the $\pi_{L}$-adic topology from $B_{\mathrm{inf},L}^{+}$ as well as the $\ker(\theta)$-adic one, along with an action of $G_{L}$ which is compatible with these topologies. Again, $\theta$ extends to a $G_{L}$-equivariant epimorphism $\theta\colon B_{\mathrm{dR}}^{+}\twoheadrightarrow\mathbb{C}_{p}$ and its kernel remains principal. The $\mathbb{Z}$-powers of the ideal $\ker(\theta)$ induce a decreasing, exhaustive and separated filtration $(\operatorname{Fil}^{n}B_{\mathrm{dR}})_{n\in\mathbb{Z}}$ on the fraction field $B_{\mathrm{dR}}$ of $B_{\mathrm{dR}}^{+}$. This field is our first ring of interest.
The field $B_{\mathrm{dR}}$ is sufficient to capture $p$-adic representations arising from formal groups, but too coarse to be used to classify them. Let $\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}(G_{L})$ be the category of $p$-adic representations of $G_{L}$, and let $\operatorname{\mathbf{Fil}}_{L}$ be the category of finite-dimensional $L$-vector spaces equipped with a decreasing, exhaustive and separated $\mathbb{Z}$-filtration by subspaces. Fontaine defined a contravariant functor
@@
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{dR}})$, equipped with the induced filtration $\operatorname{Fil}^{n}\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,\operatorname{Fil}^{n}B_{\mathrm{dR}})$ for all $n\in\mathbb{Z}$. We say that a $p$-adic representation $V$ of $G_{L}$ is de Rham if the $L$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. In this case, the jumps of the filtration are the Hodge-Tate weights of $V$. Restricting the latter functor to de Rham representations yields a faithful functor. When $(V,\rho)$ is one-dimensional, being de Rham is equivalent to the existence of a nonzero period $z\in B_{\mathrm{dR}}$ such that $g.z=\rho(g)z$ for all $g\in G_{L}$. Then its weight is the largest $w\in\mathbb{Z}$ such that $z\in\operatorname{Fil}^{w}B_{\mathrm{dR}}$.
@@ 2.2.4. The crystalline period ring
To refine the theory, Fontaine defined a subring $B_{\mathrm{cris}}$ of $B_{\mathrm{dR}}$ as follows: let $A_{\mathrm{cris}}^{\circ}$ be the PD-envelope of $\operatorname{W}(\mathcal{R})$ with respect to $\ker(\theta)$. Set $A_{\mathrm{cris}}$ to be its $p$-adic completion; one can show that it is still a subring of $B_{\mathrm{dR}}$. Define $B_{\mathrm{cris}}^{+}=A_{\mathrm{cris}}[1/p]$. The $G_{L}$-action on $\operatorname{W}(\mathcal{R})$ extends at each previous ring and the latter contains a unique $G_{L}$-stable $\mathbb{Q}_{p}$-line on which the group $G_{L}$ acts by the cyclotomic character. The crystalline period ring
$B_{\mathrm{cris}}=B_{\mathrm{cris}}^{+}[1/t]$
is obtained from $B_{\mathrm{cris}}^{+}$ by inverting any nonzero $t$ in this line. Unlike in the case of $B_{\mathrm{dR}}$, the Frobenius $\varphi$ on $\operatorname{W}(\mathcal{R})$ extends uniquely to $B_{\mathrm{cris}}$, where $\varphi(t)=pt$, into an injective $G_{L}$-equivariant endomorphism $\varphi\colon B_{\mathrm{cris}}\hookrightarrow B_{\mathrm{cris}}$ which is semilinear with respect to the Frobenius $\sigma$ of $L_{0}$, the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$. The canonical map $L\otimes_{L_{0}}B_{\mathrm{cris}}\rightarrow B_{\mathrm{dR}}$ is injective and $G_{L}$-equivariant.
The ring $B_{\mathrm{cris}}$ will be of primary interest here, as it is rich enough not only to capture but also to classify representations arising from formal groups. Let $\operatorname{\mathbf{MF}}_{L}(\varphi)$ be the category of finite-dimensional $L_{0}$-vector spaces with a bijective $\sigma$-semilinear Frobenius endomorphism $\varphi$ and whose scalar extension to $L$ is an object of $\operatorname{\mathbf{Fil}}_{L}$. Morphisms in this category are $L_{0}$-linear maps commuting with $\varphi$ and preserving the filtration after scalar extension. Fontaine defined a contravariant functor
@@
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{cris}})$, equipped with the Frobenius induced by that on $B_{\mathrm{cris}}$. One has $L\otimes_{L_{0}}\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)\hookrightarrow\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$. We say that a $p$-adic representation $V$ of $G_{L}$ is crystalline if the $L_{0}$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. The Hodge-Tate weights of a crystalline representation are its weights as a de Rham representation. Restricting the previous functor to crystalline representations yields an anti-equivalence of categories
@@ 2.2.5. Application to formal groups
Given a formal group $H$ of height $h\geqslant 1$ defined over $\mathcal{O}_{L}$, the $\mathbb{Q}_{p}$-scalar extension of its $p$-adic Tate module
$\operatorname{V}_{\!p}(H)=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\operatorname{T}_{p}(H)$
is a de Rham $G_{L}$-representation of dimension $h$ and of Hodge-Tate weights 0 with multiplicity h1h-1 and $1$ with multiplicity $1$. We say that such a representation is of formal group type. Moreover, $\operatorname{V}_{\!p}(H)$ is a crystalline representation of $G_{L}$. In this paper, we are primarily interested in the integral setting for which the situation is as follows. Denote by $\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$ the category of formal groups of finite height defined over $\mathcal{O}_{L}$ and by $\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},\mathrm{fgl}}(G_{L})$ the category of $G_{L}$-stable $\mathbb{Z}_{p}$-lattices in crystalline $G_{L}$-representations of formal group type. According to [SW13, Corollary 6.2.3] and [Tat67, Proposition 1], the functor
$\operatorname{T}_{p}\colon\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}\longrightarrow\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},\mathrm{fgl}}(G_{L})$
is an equivalence. When $H$ is of height $1$, any endomorphism $s\in\operatorname{End}_{\mathcal{O}_{L}}(H)$ acts on $\operatorname{T}_{p}(H)$ by multiplication by its linear coefficient $s^{\prime}(0)\in\mathbb{Z}_{p}$, and hence extends to an endomorphism of $\operatorname{V}_{\!p}(H)$. In view of this equivalence and the fact that this scalar extension is faithful, the single coefficient $s^{\prime}(0)$ together with the $\mathbb{Z}_{p}$-module structure on $\operatorname{T}_{p}(H)$ or $\operatorname{V}_{\!p}(H)$ uniquely determine the endomorphism $s$. It also acts on $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(\operatorname{V}_{\!p}(H))$ in the following way: for a nonzero $v\in\operatorname{V}_{\!p}(H)$ and for a nonzero period $z\in B_{\mathrm{dR}}$ of $\operatorname{V}_{\!p}(H)$, the map induced by $s$ sends $(v\mapsto z)$ to $(v\mapsto s^{\prime}(0)z)$ since $s(v)=s^{\prime}(0)v$. The initial series $s$ is unequivocally determined by this action.